What is the pH at the equivalence point for the titration of 0.23 M solutions of the following acids and bases.
NaHSO3 (for H2SO3: Ka1=1.5×10−2, Ka2=6.3×10−8) and NaOH
It says my answer 4.08 is incorrect.
1
Expert's answer
2018-07-30T06:06:27-0400
Moles NaHSO3 = 0.23
Moles NaOH required = 0.23
Volume NaOH = 0.23mol/0.23M = 1L
Total volume = 2 L
HSO3^- + OH^- = SO3^2- + H2O
[SO3^2-]=0.23mol/2L = 0.115 M
SO3^2- + H2O = HSO3^- + OH^-
6.3×10^-8 = x^2/0.115-x
x = [H3O^+]= 8.5×10^-5
pH = 4.07